Correctly answer each part of question 6
Ratios are mcstly calculated using data drawn from the financial statements of a firm. However, another group of ratios, called market-based ratios, relate to a firms observable market value, stock prices, and book values integrating information from both the market and the firms financial statements Consider the case of Cute Camel Woodcraft Company: Cute Camel Woodcraft Company just reported earnings after tax (also called net income) of $9,750,000, and a current stock price of $23.00 per share. The company is forecasting an increase of 25% for its after-tax income next year, but it also expects it will have to issue 2,900,000 new shares of stock (raising its shares outstanding from 5,500,000 to 8,400,000). If Cute Camels forecast turns out to be correct and its price-to-earnings (P/E) ratio does not change, what does the companys management expect its stock price to be one year from now? (Round any P/E ratio calculation to four decimal places). $18.84 per share $23.00 per share $14.13 per share $23.55 per share One year later, Cute Camels shares are trading at $52.08 per share, and the company reports the value of its total common equity as $31,651,200. Given this information, Cute Camels market-to-book (M/B) ratio is
One year later, Cute Camels shares are trading at $52.08 per share, and the company reports the value of its total common equity as $31,651,200. Given this information, Cute Camels market-to-book (M/B) ratio is Is it possible for a company to exhibit a negative EPS and thus a negative P/E ratio? O Yes O No Which of the following statements is true about market value ratios? E ratios could mean that the company has a great deal of uncertainty in its future earnings O Low P/E ratios could mean that the company has a great deal of uncertainty in its future earnings. Fast, Pierer w1% 32.С,0,114 Q3 3.341 2004-2016 Api.. An nghts reserved o 201 Cengaga Learvng exnopt as neted: All rights reserved Grade it Now Save & Continue
Answer #1

Question 1:

Current Net income = $ 9750000

Current earnings per share = Current Net income/no of shares = $ 9750000/5500000=$ 1.77

Current Stock price = $ 23 per share

Current PE ratio = Current Stock price per share/ Current earnings per share =23/1.77=12.97

New net income = (1+25%)*$ 9750000=$12187500

New earnings per share = New Net income/no of shares = $12187500/8400000=$ 1.45

PE ratio is the same as per data in the question = 12.97

Therefore new stock price per share = New earnings per share* PE ratio = $ 1.45*12.97=$18.84

Question 2:

Market to book value = market value of share / book value = (price of share * no of shares) / book value = $ (52.08*8400000)/ $31651200=13.82

Question 3:

This is technically possible but its generally not reported at all (if value is negative)

Question 4:

If PE ratio is low, it means the market is currently placing lower premium for every $ of earnings and the reason could be that there could be some / perceived uncertainty of its future earnings and vice versa